OpenStudy (anonymous):

What is the slope of the line 5x – 2y = -10?

OpenStudy (anonymous):

slope from standard form is -a/b so, -5/2

OpenStudy (anonymous):

first make everthing equal to y so.... -2y=-5x-10 y=5/2x+5 then you just look at the number by the x which is 5/2 and that would be your slope

OpenStudy (anonymous):

or that. but im still convinced that its -a/b

OpenStudy (amistre64):

no need to add extra work into it unless you really like to ;) ax+by=c ; slope = -a/b

OpenStudy (anonymous):

so therefore -5/2

OpenStudy (amistre64):

-5/-2 = 5/2

OpenStudy (anonymous):

thank u amistre

OpenStudy (anonymous):

oh, whoops, nvrmind, missed th -2 :D

OpenStudy (amistre64):

ax+by = c by = -ax + c y = (-a/b)x + c/b

OpenStudy (anonymous):

there is a negative people! the answer is most definately positive! :)

OpenStudy (anonymous):

ya, ik, sorry